Jean is 54 and Kay is 2 years old. When will Jean be 3 times Kay's age? Please show me how to write out the problem. I am lost.
Answer by ewatrrr(24785)

x the number of years that past
54 + x = 3(2 + x)
54 + x = 6 + 3x
48 = 2x
24 = x, In 24 years
And...
54 + 24 = 3*26